OpenStudy (anonymous):

Compare and contrast, in complete sentences, the difference between the products of a2a4 and (a2)4. Provide the expanded form and product of each.

OpenStudy (cwrw238):

when terms with exponents are multiplied the exponents are added. whereas when a term with exponents is itself subject to a further exponent the exponents are multiplied

OpenStudy (cwrw238):

so a^2 * a^4 = a^(2+4) and (a^2)^4 = a ^(2*4)

OpenStudy (anonymous):

\[\huge x^a \times x^b = x^{a+b}\] \[\huge (x^a)^b = x^{ab}\]